OpenStudy (jazzyfa30):

help please z^5-3z^2-20/z-2

OpenStudy (anonymous):

So what do they want you to do?

OpenStudy (anonymous):

Help with what

OpenStudy (anonymous):

@jazzyfa30

OpenStudy (jazzyfa30):

simplify

OpenStudy (anonymous):

How would you like the rational function to be simplified? As the ratio of two polynomials?

OpenStudy (anonymous):

Ok so if you use synthetic division it is pretty easy...just remember that any terms that are not there must be represented by 0's

OpenStudy (anonymous):

The term simplified is vague with out any context, in different situations different expressions might be seen as simpler as one or another.

OpenStudy (jazzyfa30):

ummmmm ok and it dont matter @jack117

OpenStudy (anonymous):

You want it in the form $$\frac{Q(z)}{P(z)}$$ Where Q, and P are polynomials with integer coeiffients?

OpenStudy (anonymous):

I assume the original expression is \[\frac{z^5-3z^2-20}{z-2}\]

OpenStudy (anonymous):

Most symbolic calculators can do this sort of thing, with out much trouble so in general there isn't really need to post for help here if you have a problem like this, you could try somthing free like wolframalpha.

OpenStudy (anonymous):

Do you know how to do synthetic division @jazzyfa30

OpenStudy (jazzyfa30):

yes paynesdad

OpenStudy (jazzyfa30):

??????/////

OpenStudy (anonymous):

Ok so what would be your line of coefficients?

OpenStudy (anonymous):

\[Z^5-3z^2-20/(z-2)\] multiply by the inverse of 1/(z-2) \[=z^5-3z^2-20z+40\]

OpenStudy (anonymous):

@jazzyfa30

OpenStudy (anonymous):

@NickChio6 not quite

OpenStudy (anonymous):

If the numerator is the first polynomial, then\[\frac{z^5-3z^2-20}{z-2}\]is a division problem that can be done by long division or synthetic division.

OpenStudy (jazzyfa30):

ok im starting to get confused

OpenStudy (anonymous):

\[=z(z^4-3z-20)+40\]

OpenStudy (anonymous):

Ok WIth synthetic division you list all of the coefficients of the thing you are dividing... right?

OpenStudy (anonymous):

@NickChio6 no

OpenStudy (jazzyfa30):

idk

OpenStudy (anonymous):

For example ...\[\frac{x^3-x^2-x-15}{x-3}\]

OpenStudy (anonymous):

well it depends on if it is all divisible by z-2 or just -20 is like it is written.

OpenStudy (anonymous):

You would do list the coefficientse of the numerator... 1 -1 -1 -15

OpenStudy (anonymous):

Does that look familiar at all?

OpenStudy (anonymous):

@NickChio6 It is divisible by z-2 you just need to do the synthetic division

OpenStudy (anonymous):

Look at what I asked her above... as a clarification of the problem.

OpenStudy (anonymous):

@jazzyfa30

OpenStudy (jazzyfa30):

nope

OpenStudy (anonymous):

So do you know how to do synthetic division or not?

OpenStudy (anonymous):

What course is this for?

OpenStudy (jazzyfa30):

algebra 2

OpenStudy (anonymous):

So you should have learned or should be learning synthetic division.

OpenStudy (jazzyfa30):

nvm i got it thanks anyway

OpenStudy (anonymous):

How did you do the problem?

OpenStudy (jazzyfa30):

ohh i just guessed

OpenStudy (anonymous):

:-( That is not going to help you learn how to do this stuff.